# Six decades of the FitzHugh-Nagumo model: A guide through its spatio-temporal dynamics and influence across disciplines

### v2

@@ I Introduction
Mathematical modeling is a powerful tool used in various fields to study and understand complex phenomena Meerschaert (2013). It involves the use of mathematical equations and techniques to represent real-world problems, systems, or processes in a simplified and abstract form. This allows analyzing and predicting the behavior of the system under different conditions, and make informed decisions based on the results of the analysis. Mathematical modeling is widely used in physics Redhead (1980), engineering Hritonenko and Yatsenko (2003), finance Elliott and Kopp (2005), biology Edelstein-Keshet (2005), medicine Ottesen and Danielsen (2000), and many other fields Pabjan (2004); Henley (1976). It has led to significant advances in scientific research, technological innovation, and practical problem-solving. In this way, mathematical modeling has become an indispensable tool in modern science and society.
In this work, we review the FitzHugh-Nagumo (FHN) equation as one successful example of mathematical modeling (Fig. 1D). We will briefly discuss how the FHN model was introduced in the field of neuroscience, becoming a widely used model for neuronal dynamics. However, the FHN model exists in many variations of the original system, and it has been studied in many different fields due to its simplicity and its rich and generic dynamical behavior, which we highlight here. The original FHN equation shows bistability, (relaxation) oscillations and excitability. When coupling multiple FHN models, for example through spatial diffusive coupling, traveling waves and extended patterns can be found. In this review, we will highlight many of the most commonly found dynamical states and bifurcations.
@@ I.1 Action potentials in neurons: the Hodgkin-Huxley model
Information in nerve fibers is encoded through electrical membrane changes, known as action potentials. In 1939, Alan Hodgkin and Andrew Huxley managed to make the first intracellular recording of an action potential by inserting microelectrodes into the giant axons of squids (Fig. 1A). In their experiments, Hodgkin and Huxley were able to demonstrate that this action potential was the result of two distinct contributions: a rapid inward current carried by sodium (Na+) ions, and a more slowly activating outward current carried by potassium (K+) ions. At rest, a cell is not actively transmitting a signal, and there is a potential difference of about 70 mV across the membrane. When a sufficiently strong external stimulus disrupts the cell, the membrane potential rapidly increases, leading to a phenomenon known as depolarization. Subsequently, the membrane potential decreases back to its baseline level, a phase referred to as repolarization.
Using voltage-clamp methods, they discovered that the permeability of the membrane for sodium and potassium was regulated independently, with the conductances depending on both time and the membrane potential. Remarkably, they provided a complete computational model for the action potential in a single cell, now known as the Hodgkin-Huxley (HH) model Hodgkin and Huxley (1952), which was a major advancement of our understanding. For their seminal work, they were awarded the Nobel Prize in Physiology or Medicine in 1963 Nob .
@@
$\begin{array}[]{ @{} l @{} }I=C_{m}\frac{dV_{m}}{dt}+\overline{g}_{K}n^{4}(V_{m}-V_{k})+\\ \\ \overline{g}_{Na}m^{3}h(V_{m}-V_{Na})+\overline{g}_{l}(V_{m}-V_{l})\\ \\[3.0pt] \displaystyle\frac{dn}{dt}=\alpha_{n}(V_{m})(1-n)-\beta_{n}(V_{m})n\\ \\[3.0pt] \displaystyle\frac{dm}{dt}=\alpha_{m}(C_{m})(1-m)-\beta_{m}(V_{m})m\\ \\[3.0pt] \displaystyle\frac{dh}{dt}=\alpha_{h}(V_{m})(1-h)-\beta_{h}(V_{m})h\end{array},$ (1)
@@
The fact that the HH model correctly captures these excitable dynamics Hodgkin and Huxley (1952) was a major realization as it provided a detailed understanding of the measured action potential and the dynamical process of neural excitability (Fig. 1C). E: FitzHugh (1961); Izhikevich (2007a); Izhikevich and FitzHugh (2006); Upadhyay and Iyengar (2021); Massanés and Vicente (1999); Sherwood (2014); Siettos and Starke (2016); Deco et al. (2008) TW: Nagumo et al. (1962); Izhikevich and FitzHugh (2006); Bressloff (2014); Sherwood (2014) S: Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020); Chun-Biao et al. (2010); Al Beattie et al. (2024) O: FitzHugh (1961); Izhikevich (2007a); Izhikevich and FitzHugh (2006); Upadhyay and Iyengar (2021); Sherwood (2014); Longtin (1993) SP: Murray (2001) O: Goetze and Lai (2022) SB: FitzHugh (1961); Izhikevich (2007a); Upadhyay and Iyengar (2021) E: Goetze and Lai (2022) E: Clayton and Panfilov (2008); Keener and Sneyd (2009) SW: Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Van Capelle and Durrer (1980); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008) TW: Keener (1988) O: Keener (1988) TW: Franzone et al. (2006); Keener (1988); Keener and Sneyd (2009); Biktashev (2002); Glass and Josephson (1995); Nagai et al. (2000) OP: Watanabe et al. (2004); Le and Sotiropoulos (2013); Panfilov and Keener (1995) O: Ullner et al. (2003); Rosin et al. (2011); Romeira et al. (2016); Linares-Barranco et al. (1991); Keener (1983); Khakipoor et al. (2022); Xu et al. (2020); Liu et al. (2020); Barland et al. (2003); Li et al. (2021); Binczak et al. (2006, 2003) S: Ciszak et al. (2003); Zhang and Liao (2017); Li et al. (2021); Binczak et al. (2006) SB: Romeira et al. (2016); Xu et al. (2020); Bao et al. (2019); Zhang and Liao (2017); Liu et al. (2020); Barland et al. (2003); Binczak et al. (2006, 2003) SB: Heinrich et al. (2010); Zhang and Liao (2017); Li et al. (2021) E Rosin et al. (2011); Romeira et al. (2016); Liu et al. (2020); Barland et al. (2003); Binczak et al. (2006, 2003) O: Li et al. (2021); Takembo et al. (2019) SP: Takembo et al. (2019) OP: Takembo et al. (2019) O: Ginoux and Llibre (2015); Rotstein et al. (2012); Lindner (2004); Wiesenfeld et al. (1994); Desroches et al. (2012, 2008); Makarov et al. (2001); Pankratova et al. (2005); Baer and Erneux (1992, 1986); Kostova et al. (2004); Prokop and Gelens (2024); Prokop et al. (2024) TW: McKean Jr (1970); Liu and Van Vleck (2006); Champneys et al. (2007); Majumdar and Lahiri (2007); Yochelis et al. (2008); Zemskov et al. (2008); Zemskov and Loskutov (2008); Ei et al. (2008); Stich et al. (2009); Zemskov et al. (2015); Rinzel and Keller (1973); Yochelis et al. (2015); Ziepke et al. (2016); Tyson and Keener (1988a); Elphick et al. (1997); Krupa et al. (1997); Bär et al. (2000); Ikeda et al. (2000); Rabinovitch et al. (2001); Elphick et al. (1990); Hagberg and Meron (1994a); Tsyganov et al. (2014); Rombouts and Gelens (2020); Zaritski et al. (2004); Shepelev et al. (2017); Browne et al. (2008); Guckenheimer and Kuehn (2010, 2009); Berezovskaya et al. (2008); Zemskov et al. (2011, 2018); Zemskov and Epstein (2010); Hastings (1982); Bode (1997); Tobias and Knobloch (1998); Rombouts and Gelens (2021) S: Vaidyanathan (2015); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Lindner (2004); Neiman et al. (1999); Schöll (2016a); Omelchenko et al. (2013); Semenova et al. (2016); Schöll (2016b); Omelchenko et al. (2015); Perc (2007, 2005a); Steur et al. (2009); Zaks et al. (2005); Schmidt et al. (2017); Gerster et al. (2020); Yanagita et al. (2005); Wang et al. (2009); Toral et al. (2003b); Yu et al. (2021); Brandstetter et al. (2010); Nikitin et al. (2019); Sawicki et al. (2019); Panchuk et al. (2013); Schülen et al. (2019); Schöll et al. (2009); Perlikowski et al. (2010); Saha and Feudel (2017) SB: Rocsoreanu et al. (2012); Lindner (2004); Baer and Erneux (1992, 1986); Kostova et al. (2004) SP: Bär et al. (2000); Ohta et al. (1989); Iqbal et al. (2017); Meron (1992); Petrich and Goldstein (1994); Goldstein et al. (1996); Vasiev (2004); Zheng and Shen (2015); Hagberg et al. (2006); Hagberg and Meron (1994b); Li et al. (2022); Kuznetsov et al. (2017); Gambino et al. (2022); Métens et al. (1997); Shoji et al. (2007) SB: Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Schülen et al. (2022); Yanagita et al. (2005); Wang et al. (2009); Burić and Todorović (2003); Zhen and Xu (2010); Achouri et al. (2022); Jia et al. (2015); Perlikowski et al. (2010); Saha and Feudel (2017) E: Lindner and Schimansky-Geier (1999); Lindner (2004) SW: Ambrosio and Aziz-Alaoui (2012); Keener (1986a); Tyson and Keener (1988a); Bär et al. (2000); Meron (1992); Panfilov and Hogeweg (1993); Winfree (1991a, b); Hagberg and Meron (1994b); Li et al. (2022); Bär and Eiswirth (1993); Shima and Kuramoto (2004); Hakim and Karma (1999); Lin et al. (2004) O: Campbell and Waite (2001); Ambrosio and Aziz-Alaoui (2012); Zaks et al. (2005); Yanagita et al. (2005); Burić and Todorović (2003); Perlikowski et al. (2010); Saha and Feudel (2017) SB: Ei et al. (2008); Iqbal et al. (2017); Zheng and Shen (2015); Hagberg et al. (2006); Hagberg and Meron (1994b); Shepelev et al. (2017); Guckenheimer and Kuehn (2010, 2009); Berezovskaya et al. (2008); Tobias and Knobloch (1998) OP: Neiman et al. (1999); Schülen et al. (2022); Perc (2007); Schmidt et al. (2017); Sawicki et al. (2019) OP: Bär et al. (2000); Meron (1992); Hagberg and Meron (1994a); Vasiev (2004); Nekhamkina and Sheintuch (2006); Zaritski et al. (2004); Perc (2005b); Lin et al. (2004) SW: Ambrosio and Aziz-Alaoui (2012); Lindner (2004); García-Ojalvo and Schimansky-Geier (1999) O: Champneys et al. (2007); Stich et al. (2009); Ambrosio and Aziz-Alaoui (2012); Rabinovitch et al. (2001); Elphick et al. (1990); Rombouts and Gelens (2020, 2021) SP: Lei and Yang (2021); Perlikowski et al. (2010) TW: Sawicki et al. (2019) O: Parra-Rivas et al. (2023a); Sneyd et al. (2017a); Hartzell et al. (2015); Tsai et al. (2008); Tyler et al. (2022) TW: Gelens et al. (2014); Noguchi (2023); Nolet et al. (2020a); Miller and Dunkel (2020); Duffy et al. (1998); Lee (1997) S: Kori et al. (2012); Cartwright (2000); Scialla et al. (2021) E: Parra-Rivas et al. (2023a); Hartzell et al. (2015) O: Gelens et al. (2014); Nolet et al. (2020a); Sneyd et al. (2017b); Lee (1997) SB: Vodeneev et al. (2011); Sukhov et al. (2011) SB: Chiangga et al. (2016); Stratford et al. (2019) SB: Gambino et al. (2023); Sneyd et al. (2017b) TW: Vodeneev et al. (2011); Sukhov et al. (2011) SP: Gambino et al. (2022, 2023) O: Kori et al. (2012) E: Noguchi (2023) SW: Lee (1997)
@@ I.2 A simplified description of neuronal excitability: the FitzHugh-Nagumo model
FitzHugh introduced the parameter $z$ to represent membrane current density, the variable $x$ is related to the membrane voltage and the sodium activation, and the variable $y$ corresponds to the sodium inactivation and the potasium activation. In this context, these equations thus work as an activator-inhibitor model. Later, Nagumo, Arimoto, and Yoshizawa proved the equivalence of this model with a electrical circuit. Fig. 2A shows this electrical equivalent, which includes a capacitor (representing membrane capacitance $C$), a tunnel diode [$F(V)$], a resistor (representing channel resistance $R$), an inductor ($L$), and a battery ($E$). The dynamics of this electrical circuit are captured by the following equations:

### v3

@@ I Introduction
Mathematical modeling is a powerful tool used in various fields to study and understand complex phenomena Meerschaert (2013). It involves the use of mathematical equations and techniques to represent real-world problems, systems, or processes in a simplified and abstract form. This allows analyzing and predicting the behavior of the system under different conditions, and making informed decisions based on the results of the analysis. Mathematical modeling is widely used in physics Redhead (1980), engineering Hritonenko and Yatsenko (2003), finance Elliott and Kopp (2005), biology Edelstein-Keshet (2005), medicine Ottesen and Danielsen (2000), and many other fields Pabjan (2004); Henley (1976). It has led to significant advances in scientific research, technological innovation, and practical problem-solving. In this way, mathematical modeling has become an indispensable tool in modern science and society.
In this work, we review the FitzHugh-Nagumo (FHN) equation as one successful example of mathematical modeling (Fig. 1D). We will briefly discuss how the FHN model was introduced in the field of neuroscience, becoming a widely used model for neuronal dynamics. However, the FHN model exists in many variations of the original system, and it has been studied in many different fields due to its simplicity and its rich and generic dynamical behavior, which we highlight here. The original FHN equation shows bistability, (relaxation) oscillations, and excitability. When coupling multiple FHN models, for example through spatial diffusive coupling, traveling waves and extended patterns can be found. In this review, we will highlight many of the most commonly found dynamical states and bifurcations.
@@ I.1 Action potentials in neurons: the Hodgkin-Huxley model
Information in nerve fibers is encoded through electrical membrane changes, known as action potentials. In 1939, Alan Hodgkin and Andrew Huxley managed to make the first intracellular recording of an action potential by inserting microelectrodes into the giant axons of squid (Fig. 1A). In their experiments, Hodgkin and Huxley were able to demonstrate that this action potential was the result of two distinct contributions: a rapid inward current carried by sodium (Na+) ions, and a more slowly activating outward current carried by potassium (K+) ions. At rest, a cell is not actively transmitting a signal, and there is a potential difference of about 70 mV across the membrane. When a sufficiently strong external stimulus disrupts the cell, the membrane potential rapidly increases, leading to a phenomenon known as depolarization. Subsequently, the membrane potential decreases back to its baseline level, a phase referred to as repolarization.
Using voltage-clamp methods, they discovered that the permeability of the membrane for sodium and potassium was regulated independently, with the conductance depending on both time and the membrane potential. Remarkably, they provided a complete computational model for the action potential in a single cell, now known as the Hodgkin-Huxley (HH) model Hodgkin and Huxley (1952), which was a major advancement of our understanding. For their seminal work, they were awarded the Nobel Prize in Physiology or Medicine in 1963 Nob .
@@
$\begin{array}[]{ @{} l @{} }I=C_{m}\frac{dV_{m}}{dt}+\overline{g}_{K}n^{4}(V_{m}-V_{k})+\\ \\ \overline{g}_{Na}m^{3}h(V_{m}-V_{Na})+\overline{g}_{l}(V_{m}-V_{l})\\ \\[3.0pt] \displaystyle\frac{dn}{dt}=\alpha_{n}(V_{m})(1-n)-\beta_{n}(V_{m})n\\ \\[3.0pt] \displaystyle\frac{dm}{dt}=\alpha_{m}(V_{m})(1-m)-\beta_{m}(V_{m})m\\ \\[3.0pt] \displaystyle\frac{dh}{dt}=\alpha_{h}(V_{m})(1-h)-\beta_{h}(V_{m})h\end{array},$ (1)
@@
The fact that the HH model correctly captures these excitable dynamics Hodgkin and Huxley (1952) was a major realization as it provided a detailed understanding of the measured action potential and the dynamical process of neural excitability (Fig. 1C). E: FitzHugh (1961); Izhikevich (2007a); Izhikevich and FitzHugh (2006); Upadhyay and Iyengar (2021); Massanés and Vicente (1999); Sherwood (2014); Siettos and Starke (2016); Deco et al. (2008) TW: Nagumo et al. (1962); Izhikevich and FitzHugh (2006); Bressloff (2014); Sherwood (2014) S: Goetze and Lai (2022); Brown et al. (2004); Stefanescu and Jirsa (2008); Chouzouris et al. (2018); Gerster et al. (2020); Chun-Biao et al. (2010); Al Beattie et al. (2024) O: FitzHugh (1961); Izhikevich (2007a); Izhikevich and FitzHugh (2006); Upadhyay and Iyengar (2021); Sherwood (2014); Longtin (1993) SP: Murray (2001) O: Goetze and Lai (2022) SB: FitzHugh (1961); Izhikevich (2007a); Upadhyay and Iyengar (2021) E: Goetze and Lai (2022) E: Clayton and Panfilov (2008); Keener and Sneyd (2009) SW: Davidenko et al. (1995); Courtemanche et al. (1990); Aliev and Panfilov (1996); Van Capelle and Durrer (1980); Pertsov et al. (1993); Nash and Panfilov (2004); Sermesant et al. (2006); Cherubini et al. (2008); Karma (1993); Alonso et al. (2003, 2016) TW: Keener (1988) O: Keener (1988) TW: Franzone et al. (2006); Keener (1988); Keener and Sneyd (2009); Biktashev (2002); Glass and Josephson (1995); Nagai et al. (2000) OP: Watanabe et al. (2004); Le and Sotiropoulos (2013); Panfilov and Keener (1995) O: Ullner et al. (2003a); Rosin et al. (2011); Romeira et al. (2016); Linares-Barranco et al. (1991); Keener (1983); Khakipoor et al. (2022); Xu et al. (2020); Liu et al. (2020); Barland et al. (2003); Li et al. (2021); Binczak et al. (2006, 2003) S: Ciszak et al. (2003); Zhang and Liao (2017); Li et al. (2021); Binczak et al. (2006) SB: Romeira et al. (2016); Xu et al. (2020); Bao et al. (2019); Zhang and Liao (2017); Liu et al. (2020); Barland et al. (2003); Binczak et al. (2006, 2003) SB: Heinrich et al. (2010); Zhang and Liao (2017); Li et al. (2021) E Rosin et al. (2011); Romeira et al. (2016); Liu et al. (2020); Barland et al. (2003); Binczak et al. (2006, 2003) O: Li et al. (2021); Takembo et al. (2019) SP: Takembo et al. (2019) OP: Takembo et al. (2019) O: Ginoux and Llibre (2015); Rotstein et al. (2012); Lindner (2004); Wiesenfeld et al. (1994); Desroches et al. (2012, 2008); Makarov et al. (2001); Pankratova et al. (2005); Baer and Erneux (1992, 1986); Kostova et al. (2004); Prokop and Gelens (2024); Prokop et al. (2024) TW: McKean Jr (1970); Liu and Van Vleck (2006); Champneys et al. (2007); Majumdar and Lahiri (2007); Yochelis et al. (2008); Zemskov et al. (2008); Zemskov and Loskutov (2008); Ei et al. (2008); Stich et al. (2009); Zemskov et al. (2015); Rinzel and Keller (1973); Yochelis et al. (2015); Ziepke et al. (2016); Tyson and Keener (1988a); Elphick et al. (1997); Krupa et al. (1997); Bär et al. (2000); Ikeda et al. (2000); Rabinovitch et al. (2001); Elphick et al. (1990); Hagberg and Meron (1994a); Tsyganov et al. (2014); Rombouts and Gelens (2020); Zaritski et al. (2004); Shepelev et al. (2017); Browne et al. (2008); Guckenheimer and Kuehn (2010, 2009); Berezovskaya et al. (2008); Zemskov et al. (2011, 2018); Zemskov and Epstein (2010); Hastings (1982); Bode (1997); Tobias and Knobloch (1998); Rombouts and Gelens (2021); Panfilov and Rudenko (1987); Saha et al. (2020); García-Ojalvo et al. (2001); Brauns and Marchetti (2024) S: Vaidyanathan (2015); Toral et al. (2003a); Ambrosio and Aziz-Alaoui (2012); Lindner (2004); Neiman et al. (1999); Schöll (2016a); Omelchenko et al. (2013); Semenova et al. (2016); Schöll (2016b); Omelchenko et al. (2015); Perc (2007, 2005a); Steur et al. (2009); Zaks et al. (2005); Schmidt et al. (2017); Gerster et al. (2020); Yanagita et al. (2005); Wang et al. (2009); Toral et al. (2003b); Yu et al. (2021); Brandstetter et al. (2010); Nikitin et al. (2019); Sawicki et al. (2019); Panchuk et al. (2013); Schülen et al. (2019); Schöll et al. (2009); Perlikowski et al. (2010); Saha and Feudel (2017) SB: Rocsoreanu et al. (2012); Lindner (2004); Baer and Erneux (1992, 1986); Kostova et al. (2004) SP: Bär et al. (2000); Ohta et al. (1989); Iqbal et al. (2017); Meron (1992); Petrich and Goldstein (1994); Goldstein et al. (1996); Vasiev (2004); Zheng and Shen (2015); Hagberg et al. (2006); Hagberg and Meron (1994b); Li et al. (2022); Kuznetsov et al. (2017); Gambino et al. (2022); Métens et al. (1997); Shoji et al. (2007) SB: Hoff et al. (2014); Brandibur and Kaslik (2022); Campbell and Waite (2001); Schülen et al. (2022); Yanagita et al. (2005); Wang et al. (2009); Burić and Todorović (2003); Zhen and Xu (2010); Achouri et al. (2022); Jia et al. (2015); Perlikowski et al. (2010); Saha and Feudel (2017) E: Lindner and Schimansky-Geier (1999); Lindner (2004) SW: Ambrosio and Aziz-Alaoui (2012); Keener (1986a); Tyson and Keener (1988a); Bär et al. (2000); Meron (1992); Panfilov and Hogeweg (1993); Winfree (1991a, b); Hagberg and Meron (1994b); Li et al. (2022); Bär and Eiswirth (1993a); Shima and Kuramoto (2004); Hakim and Karma (1999); Lin et al. (2004); Bär and Eiswirth (1993b); Dowle et al. (1997); Barkley (1991) O: Campbell and Waite (2001); Ambrosio and Aziz-Alaoui (2012); Zaks et al. (2005); Yanagita et al. (2005); Burić and Todorović (2003); Perlikowski et al. (2010); Saha and Feudel (2017); Ullner et al. (2003b) SB: Ei et al. (2008); Iqbal et al. (2017); Zheng and Shen (2015); Hagberg et al. (2006); Hagberg and Meron (1994b); Shepelev et al. (2017); Guckenheimer and Kuehn (2010, 2009); Berezovskaya et al. (2008); Tobias and Knobloch (1998); Brauns et al. (2020); Frohoff-Hülsmann and Thiele (2023) OP: Neiman et al. (1999); Schülen et al. (2022); Perc (2007); Schmidt et al. (2017); Sawicki et al. (2019) OP: Bär et al. (2000); Meron (1992); Hagberg and Meron (1994a); Vasiev (2004); Nekhamkina and Sheintuch (2006); Zaritski et al. (2004); Perc (2005b); Lin et al. (2004); Sagués et al. (2007); Brauns and Marchetti (2024) SW: Ambrosio and Aziz-Alaoui (2012); Lindner (2004); García-Ojalvo and Schimansky-Geier (1999) O: Champneys et al. (2007); Stich et al. (2009); Ambrosio and Aziz-Alaoui (2012); Rabinovitch et al. (2001); Elphick et al. (1990); Rombouts and Gelens (2020, 2021) SP: Lei and Yang (2021); Perlikowski et al. (2010) TW: Sawicki et al. (2019) O: Parra-Rivas et al. (2023a); Sneyd et al. (2017a); Hartzell et al. (2015); Tsai et al. (2008); Tyler et al. (2022) TW: Gelens et al. (2014); Noguchi (2023); Nolet et al. (2020a); Miller and Dunkel (2020); Duffy et al. (1998); Lee (1997) S: Kori et al. (2012); Cartwright (2000); Scialla et al. (2021) E: Parra-Rivas et al. (2023a); Hartzell et al. (2015) O: Gelens et al. (2014); Nolet et al. (2020a); Sneyd et al. (2017b); Lee (1997) SB: Vodeneev et al. (2011); Sukhov et al. (2011) SB: Chiangga et al. (2016); Stratford et al. (2019) SB: Gambino et al. (2023); Sneyd et al. (2017b) TW: Vodeneev et al. (2011); Sukhov et al. (2011) SP: Gambino et al. (2022, 2023) O: Kori et al. (2012) E: Noguchi (2023) SW: Lee (1997)
@@ I.2 A simplified description of neuronal excitability: the FitzHugh-Nagumo model
FitzHugh introduced the parameter $z$ to represent membrane current density, the variable $x$ is related to the membrane voltage and the sodium activation, and the variable $y$ corresponds to the sodium inactivation and the potassium activation. In this context, these equations thus work as an activator-inhibitor model. Later, Nagumo, Arimoto, and Yoshizawa proved the equivalence of this model with a electrical circuit. Fig. 2A shows this electrical equivalent, which includes a capacitor (representing membrane capacitance $C$), a tunnel diode [$F(V)$], a resistor (representing channel resistance $R$), an inductor ($L$), and a battery ($E$). The dynamics of this electrical circuit are captured by the following equations: